# Properties of Lyapunov Subcenter Manifolds in Conservative Mechanical Systems

### v2

@@ 3.1 Conditions on inertia tensor and potential energy
The proof relies on showing that the conditions of Theorem 4 are equivalent to those of Theorem 1. We relegate the technical details to this section of the supplementary material [34].
@@ 4.2 Time and spatially symmetric Lyapunov subcenter manifolds
$\bar{Q}$
$\left(Q,0\right)$
$\left(\varphi(Q),0\right)$
$\left(Q(t),P(t)\right)$
@@ A.1.5 Generator Property
We first show that $\mathcal{R}=\{(x,0)\,|\,(x,0)\in\mathcal{M}\}$ is a regular 1-dimensional submanifold of $\mathcal{M}$. We use the regular level set theorem [38, Corr. 5.24], which states that $\mathcal{R}\subset\mathcal{M}$ is a closed, embedded submanifold of dimension 1 if there is a function $\alpha:\mathcal{M}\rightarrow\mathbb{R}$ such that $\mathcal{R}=\alpha^{-1}(0)$ and $\text{Im}(\text{d}\alpha)=T_{0}\mathbb{R}$. To this end, note that
@@ A.2.1 Equivariant coordinates
It remains to be shown that $Y$ is a valid chart-map, i.e., that $Y$ is a local diffeomorphism. This follows from the inverse function theorem [39, Theorem 1.3.12] by showing that $Y_{*}(\bar{x})$ is full rank:
@@ A.2.2 Proof of spatial symmetry
Both sides are the same, so the symmetry condition 58 holds. This completes the proof.

### v3

# arXiv:2505.13064v3[math.DS] 02 Mar 2026

@@ 1 Introduction
1.1 Related literature
Exciting and exploiting efficient nonlinear oscillations is an open challenge in nonlinear control theory [35, 36, 37], with important ramifications in robotics’ locomotion and execution of industrial operations [38, 39, 40, 41, 42], and biomechanics [43, 44, 45]. To this end, Eigenmanifolds have been recently introduced (see [22, Sec. 7]) as an extension of modal analysis to multi-body systems, in a way that naturally lends itself to control applications: design of mechanical systems [46] and shaping of Eigenmanifolds via the systems potential energy [47, 16] has been explored. Eigenmanifolds of interest are stabilized in simulation [48, 49], on real systems [50], can be stabilized in an energy-efficient manner [15] and allow the swing-up of weakly actuated systems to high energies [14]. Eigenmanifolds have also been applied in classifying the accuracy of reduced order models in soft robotics [51, 52].
Eigenmanifold theory has its precursors in the study of nonlinear normal modes [53, 20, 54] for nonlinear mechanical systems with constant inertia tensors. Rosenberg modes are well-studied for conservative mechanical systems [18, 19, 55], and their symmetry properties are known by the condition that V(q)=V(q)V(q)=V(-q). Strict normal modes are a special case of Eigenmanifolds that exist under compatibility of the potential gradient with the geodesic flow [56].
Within the nonlinear dynamics literature, LSMs are long studied [23, 30, 31] and their persistence in the limit of small dissipation was shown [32], but this has not been connected to insights about Eigenmanifolds in multi-body dynamics. LSMs have also been applied to the model order reduction of partial differential equations [57].
A general treatment of discrete symmetries in dynamical systems is given by Michael Field [58]. Large classes of symmetries were also studied in the context of periodic orbits and nonlinear normal modes [59, 60]. However, these works largely treat symmetries on vector and Banach spaces, and do not directly apply to Eigenmanifold theory. An excellent overview to time-reversal symmetries in Hamiltonian systems is given by Lamb & Roberts [61], covering many properties of direct relevance to Eigenmanifold theory, such as the existence of time-symmetric LSMs. Alomair & Montaldi [62] investigate the interplay of a time-reversal and a spatial symmetry of Hamiltonian systems in the context of periodic orbits, proving the existence of a discrete number of doubly symmetric periodic orbits around symmetric equilibria. However, both of [61, 62] leave properties of such symmetric orbits and their families implicit.
@@ 3.1 Conditions on inertia tensor and potential energy
The proof relies on showing that the conditions of Theorem 4 are equivalent to those of Theorem 1.
First, express the Hamiltonian dynamics of a conservative mechanical system (2), (6) in canonical coordinates $(q,p)$ such that the equilibrium $(\bar{x},0)$ is mapped to $q=0$, $p=0$. Further write $z=(q,p)\in\mathbb{R}^{2n}$. The dynamics then take the form
$\dot{z}$ $:=f(z)=\begin{pmatrix}M(q)^{-1}p\\ -\frac{\partial}{\partial q}(p^{\top}M(q)^{-1}p)-\frac{\partial V}{\partial q}(q)\end{pmatrix}\,$, (17)
which can be rewritten as
$\dot{z}=Az+B(z)\,$, (18)
with
$A$ $=\frac{\partial f}{\partial z}\big((0,0)\big)=\begin{bmatrix}0&M(0)^{-1}\\ -\frac{\partial^{2}V}{\partial q\partial q}(0)&0\end{bmatrix}\,$, (19)
$B(z)$ $=f(z)-Az\,.$ (20)
The dynamics are of the required form (7) for Theorem 1, since $B(0)=0$, and $\frac{\partial B}{\partial z}(0)=0$. $B(z)$ is also analytic, given that $M(q),V(q)$ are analytic. This does not rely on the choice of a particular chart: the components of $A$ correspond to the components of a $(1,1)$-tensor on $T^{*}\mathcal{Q}$, given that $\text{d}V(\bar{x})=0$, which follows from $(0,0)$ being an equilibrium of the dynamics (5), (6).
The eigenvalues $\lambda$ of $A=\begin{bmatrix}0&B\\ C&0\end{bmatrix}$ are the solution to
$\det(\lambda I_{2n}-A)=\det(\lambda^{2}I_{n}-BC)=0\,$, (21)
such that $A$ is diagonalizable since the $(1,1)$-tensor $BC=-M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ is diagonalizable. Thus, condition $(i)$ of Theorem 1 is fulfilled.
Further, $\lambda^{2}=-\omega^{2}$ for some positive $\omega\in\mathbb{R}_{+}$, as $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ is a positive definite matrix collecting the components of the positive definite tensor $M(\bar{x})^{-1}\text{Hess}(V)(\bar{x})$, and eigenvalues $\omega_{j}^{2}$ of $M(0)^{-1}\frac{\partial^{2}V}{\partial q\partial q}(0)$ correspond to pairs of eigenvalues
$\lambda_{j,\pm}=\pm i\omega_{j}$ (22)
of $A$. Thus, condition $(ii)$ of Theorem 1 is fulfilled.
The eigenvalues $\lambda_{0,\pm}$ are non-resonant with the remaining 2n22n-2 eigenvalues $\lambda_{k,\pm}$ of $A$, since
$\frac{\lambda_{k}^{2}}{\lambda^{2}_{0,\pm}}\notin\mathbb{Z}\Rightarrow\frac{\lambda_{k}}{\lambda_{0,\pm}}\notin\mathbb{Z}\,$, (23)
where $\frac{\lambda_{k}^{2}}{\lambda^{2}_{0,\pm}}\notin\mathbb{Z}$ is guaranteed by assumption. Thus, also condition $(iii)$ of Theorem 1 is fulfilled.
Finally, a conserved quantity $H\big(z(t)\big)$ fulfilling condition $(iv)$ of Theorem 1 is given by the Hamiltonian $H\big(z(t)\big)=\frac{1}{2}p^{\top}M(q)^{-1}p+V(q)$ .
@@ 4.2 Time and spatially symmetric Lyapunov subcenter manifolds
$\bar{x}$
$\left(x,0\right)$
$\left(\varphi(x),0\right)$
$\left(x(t),P(t)\right)$
@@ A.1.5 Generator Property
We first show that $\mathcal{R}=\{(x,0)\,|\,(x,0)\in\mathcal{M}\}$ is a regular 1-dimensional submanifold of $\mathcal{M}$. We use the regular level set theorem [64, Corr. 5.24], which states that $\mathcal{R}\subset\mathcal{M}$ is a closed, embedded submanifold of dimension 1 if there is a function $\alpha:\mathcal{M}\rightarrow\mathbb{R}$ such that $\mathcal{R}=\alpha^{-1}(0)$ and $\text{Im}(\text{d}\alpha)=T_{0}\mathbb{R}$. To this end, note that
@@ A.2.1 Equivariant coordinates
It remains to be shown that $Y$ is a valid chart-map, i.e., that $Y$ is a local diffeomorphism. This follows from the inverse function theorem [65, Theorem 1.3.12] by showing that $Y_{*}(\bar{x})$ is full rank:
@@ A.2.2 Proof of spatial symmetry
Both sides are the same, so the symmetry condition 65 holds. This completes the proof.